Find the coefficient of X2 and x 3 in the expansion of (2-x)raised to the power of 6(


1
Expert's answer
2019-10-31T10:20:52-0400

According to binomial expansion formula, "(x+y)^n = \sum_{k=0}^n C_k^n x^k y^{n-k}", where "C_n^k = \frac{n!}{(n-k)!k!}" is the binomial coefficient.

Hence, "(2-x)^6 = \sum_{k=0}^6 C_k^6 (-x)^k 2^{6-k}" .

From the last formula, coefficients next to "x^2", "x^3" are "C_2^6 2^4 = 240" and "C_3^6 (-1)^3 2^3 = -160" respectively.